Kylie is a single mom with two dependent children, Tanner, age 7 and Olivia, age 11. She has AGI of $38,000 and paid $5,300 to qualified day care center for the two children. What amount of credit can Kylie receive for the child and dependent care credit?
Answer #1

child and dependent care credit = 5300*23% = $1219

Total expenses eligible for child and dependent care credit is upto 3000 for one dependent and 6000 for two or more dependent. here expenses are 5300 and number of qualified dependent are two and thus total expenses of 5300 are eligible for child and dependent care credit. percentage of deduction available is from 20%-35% of qualified expenses and this percentage depends on AGI. full 35% percentage is available to taxpayer with AGI of 15000 and this percentage decreases by 1% with every increase in AGI by 2000 and percentage for child and dependent care credit applicable here is 23% and thus child and dependent care credit = 5300*23%=1219